### v1

# Deep Operator Networks for Bayesian Parameter Estimation in PDEs

Abstract
We present a novel framework combining Deep Operator Networks (DeepONets) with Physics-Informed Neural Networks (PINNs) to solve partial differential equations (PDEs) and estimate their unknown parameters. By integrating data-driven learning with physical constraints, our method achieves robust and accurate solutions across diverse scenarios. Bayesian training is implemented through variational inference, allowing for comprehensive uncertainty quantification for both aleatoric and epistemic uncertainties. This ensures reliable predictions and parameter estimates even in noisy conditions or when some of the physical equations governing the problem are missing. The framework demonstrates its efficacy in solving forward and inverse problems, including the 1D unsteady heat equation and 2D reaction-diffusion equations, as well as regression tasks with sparse, noisy observations. This approach provides a computationally efficient and generalizable method for addressing uncertainty quantification in PDE surrogate modeling.
keywords:
Physics-Informed Neural Networks, Partial Differential Equations, Parameter Estimation, Machine Learning
journal: arXiv
[…]
[csulb]organization=Depratment of Computer Engineering and Computer Science, addressline= California State University, city=Long Beach, postcode=90840, state=CA, country=USA
1 Introduction
The modeling and simulation of complex physical systems governed by partial differential equations (PDEs) are critical in science and engineering. Traditional numerical methods for solving PDEs, while effective, often demand substantial computational resources and face challenges with high-dimensional or nonlinear problems [1, 2]. Neural networks have emerged as promising tools for approximating PDE solutions [3, 4, 5, 6]. However, these models typically require large datasets and lack mechanisms to embed physical laws into their frameworks.
Physics-Informed Neural Networks (PINNs) [7, 8] integrate governing physical laws, expressed as PDEs, directly into their loss functions. This approach enables PINNs to solve forward and inverse problems efficiently, even with limited data, by ensuring that predictions adhere to the underlying physics. By embedding physical laws into the learning process, PINNs overcome the reliance on extensive datasets required by traditional neural networks, making them more practical for real-world scenarios [9]. Applications of PINNs include solving fluid dynamics problems [10, 11], addressing inverse problems [12, 13], controlling dynamical systems [14, 15, 16], and performing uncertainty quantification [17].
Automatic differentiation (AD) is integral to the functionality of PINNs, enabling precise and efficient computation of PDE residuals, even for mixed and high-order derivatives. Its integration into machine learning frameworks such as PyTorch [18], Jax [19], and TensorFlow [20] has significantly advanced deep learning applications in scientific domains. Comprehensive surveys on AD and its relationship to deep learning can be found in Baydin et al. [21] and Margossian [22].
In scientific machine learning, probabilistic models are highly valued for estimating noise in observational data, addressing model errors, and accounting for incomplete knowledge of the underlying physics. Viana et al. [23] highlight the evolution of scientific modeling, emphasizing the role of physics-informed neural networks in reducing computational costs and enhancing modeling flexibility. Molnar et al. [24] demonstrate the application of PINNs within a Bayesian framework to reconstruct flow fields from projection data, showcasing their ability to incorporate physics priors and improve accuracy in noisy environments. Li et al. [25] explore the challenges of solving inverse problems with PINNs, advocating for a Bayesian approach to provide robust uncertainty quantification, particularly when observations are sparse or noisy.
Bayesian Neural Networks (BNNs) [26, 27, 28] provide a probabilistic framework for uncertainty quantification by assigning distributions to network weights. While posterior estimation over BNN parameters is challenging due to high-dimensionality and non-convexity, methods such as Hamiltonian Monte Carlo (HMC) and variational inference (VI) approximate the posterior effectively, albeit at higher computational costs than maximum likelihood approaches. Yang et al. [29] introduced Bayesian Physics-Informed Neural Networks (B-PINNs), combining BNNs and PINNs to address forward and inverse PDE problems in noisy conditions. By integrating physical equations into training, B-PINNs offer a practical solution for probabilistic PDE learning tasks.
Accurately estimating PDE parameters while quantifying uncertainty remains a significant challenge, particularly in complex systems with scarce or noisy data. The high computational demands and challenges in selecting suitable priors in Bayesian approaches restrict the practical use of black-box sampling methods in scientific machine learning. This presents an opportunity for new methodologies that effectively combine physics-guided training with computationally efficient Bayesian approaches.
In this work, we propose a novel method that extends the capabilities of PINNs by integrating Bayesian methodologies with Deep Operator Networks (DeepONets) to enhance parameter estimation in PDEs. Our approach relies on variational inference: starting with a prior distribution for the unknown PDE parameters, we apply Bayesian inference to estimate the posterior distributions based on available data and physics. The DeepOperator architecture modulates the forward prediction of the solution of the PDE with inverse parameter estimation. This method offers substantial computational advantages, providing uncertainty quantification comparable to B-PINNs but with significantly reduced computational costs.
We continue our discussion in the following steps: In section 2, we detail the methodology of Bayesian Deep Operator Networks, demonstrate their application to parameter estimation in PDEs, and discuss their advantages and shortcomings compared to existing approaches. Section 3 presents experimental results on a variety of PDEs, showcasing the effectiveness of our method in solving forward and inverse problems. Finally, we discuss the implications of our findings for future research and potential applications in various scientific and engineering domains in section 4.
2 Methodology
2.1 Problem definition
We start by considering a parametrized partial differential equation (PDE), defined over time span $t\in[t_{0},t_{f}]$ and spatial domain $x\in\Omega$:
${F}(t,x,y;\lambda)=0,\quad(t,x)\in[t_{0},t_{f}]\times\Omega.$ (1)
[…]
$\mathcal{D}=\{(t_{i},x_{i},y_{i}),~{}i=0,1,\ldots,N\},$ (2)
we are interested in training a probabilistic machine learning model that simultaneously approximates $y(t,x)$ and $\lambda$ based on the data eq. 2 and the PDE eq. 1. We will start from a Bayesian formulation of this problem and derive a corresponding loss function for training the model.
The posterior distribution over the latent parameter $\lambda$ given data $\mathcal{D}$ is expressed as
$p(\lambda\mid\mathcal{D})=\frac{p(\mathcal{D}\mid\lambda)\,p_{0}(\lambda)}{p(\mathcal{D})},$ (3)
where $p_{0}(\lambda)$ represents the prior distribution over $\lambda$, $p(\mathcal{D}\mid\lambda)$ is the likelihood of the observed data given parameter $\lambda$, and $p(\mathcal{D})$ is the marginal likelihood or evidence.
The likelihood term $p(\mathcal{D}\mid\lambda)$ is constructed from two components: the data likelihood $p(y_{i}\mid t_{i},x_{i},\lambda)$, which captures the relationship between the inputs $t_{i},x_{i}$ and outputs $y_{i}$, and the likelihood of the predicted solution satisfying the governing equations. The latter is achieved via the residuals of eq. 1.
Assuming independence across training data points, the joint likelihood is written as
$P(\mathcal{D}\mid\lambda)=\prod_{i=1}^{N}p(y_{i}\mid t_{i},x_{i},\lambda)\,p(\mathcal{R}_{\text{PDE}}(\mathcal{M}_{\theta}(x_{i},\lambda))).$ (4)
The data likelihood $p(y_{i}\mid x_{i},\lambda)$ is modeled with normal distribution:
$p(y_{i}\mid t_{i},x_{i},\lambda)=\mathcal{N}(y_{i};\mathcal{M}_{\theta}(t_{i},x_{i},\lambda),\sigma_{y}^{2}),$ (5)
where $\mathcal{M}_{\theta}(t_{i},x_{i},\lambda)$ is the model output for input $(t_{i},x_{i})$ given latent variable $\lambda$, and $\sigma_{y}^{2}$ is the observation noise variance. The log-likelihood is therefore
$\log p(y_{i}\mid t_{i},x_{i},\lambda)=-\frac{|\mathcal{M}_{\theta}(t_{i},x_{i},\lambda)-y_{i}|^{2}}{2\sigma_{y}^{2}}-\frac{1}{2}\log(2\pi\sigma_{y}^{2}).$ (6)
We are interested in models that, in addition to adherence to observation data, also produce likely predictions that satisfy the governing equations eq. 1. The likelihood of the residuals from the PDE is similarly modeled as:
$p(\mathcal{R}_{\text{PDE}}(\mathcal{M}_{\theta}(t_{i},x_{i},\lambda)))=\mathcal{N}(0,\sigma_{R}^{2}),$ (7)
where $\sigma_{R}^{2}$ is the variance of the residual noise. $\mathcal{R}_{\text{PDE}}$ can either be simplified as $|F(t_{i},x_{i},y_{i},\lambda)|$ or modified to include other constraints such as the boundary and initial conditions. The corresponding log-likelihood of the residuals is
$\log p(R_{\text{PDE}}(\mathcal{M}_{\theta}(x_{i},\lambda)))=-\frac{|\mathcal{R}_{\text{PDE}}(\mathcal{M}_{\theta}(t_{i},x_{i},\lambda))|^{2}}{2\sigma_{R}^{2}}-\frac{1}{2}\log(2\pi\sigma_{R}^{2}).$ (8)
We sample the posterior $p(\lambda\mid\mathcal{D})$ in equation eq. 3 using variational inference (VI). Considering an approximate posterior distribution $q(\lambda)$ we minimize the Kullback-Leibler (KL) divergence:
$\operatorname*{arg\,min}_{q(\lambda)}D_{KL}(q(\lambda)\|p(\lambda\mid\mathcal{D}))=\int q(\lambda)\log\frac{q(\lambda)}{p(\lambda\mid\mathcal{D})}d\lambda.$ (9)
Substituting eq. 3 for $p(\lambda\mid\mathcal{D})$ and rearranging terms, we obtain the well-known Evidence Lower Bound (ELBO) objective function:
[…]
Monte Carlo approximation of the integral in eq. 10 yields Expanding $p(\mathcal{D}\mid\lambda)$ using the joint likelihood, eq. 10 becomes
$\mathcal{L}_{\text{ELBO}}=\mathbb{E}_{q(\lambda)}\left[\sum_{i=1}^{N}\log p(y_{i}\mid x_{i},\lambda)+\log p(\mathcal{R}_{\text{PDE}}(\mathcal{M}_{\theta}(x_{i},\lambda)))\right]\\ -D_{KL}(q(\lambda)\|p_{0}(\lambda)).$ (11)
Finally, substituting the log-likelihoods eqs. 6 and 8, the loss function becomes
$\mathcal{L}=\mathbb{E}_{q(\lambda)}\Bigg{[}$ $\sum_{i=1}^{N}\left(-\frac{|\mathcal{M}_{\theta}(t_{i},x_{i},\lambda)-y_{i}|^{2}}{2\sigma_{y}^{2}}-\frac{1}{2}\log(2\pi\sigma_{y}^{2})\right)$
$-\frac{|\mathcal{R}_{\text{PDE}}(\mathcal{M}_{\theta}(t_{i},x_{i},\lambda))|^{2}}{2\sigma_{R}^{2}}-\frac{1}{2}\log(2\pi\sigma_{R}^{2})\Bigg{]}-D_{KL}(q(\lambda)\|p_{0}(\lambda)).$ (12)
Equation 12 can be evaluated using Monte Carlo sampling of the latent variable $\lambda$ from the posterior $q(\lambda)$. The observation and residual noise variances $\sigma_{y}^{2}$ and $\sigma_{R}^{2}$ can be considered constants or learned during training. This enables a flexible uncertainty quantification method that can be utilized to estimate model or data uncertainty. We experiment with different variance estimation strategies in the experiments section 3.
The KL divergence term in eq. 12 can be approximated empirically or analytically: If both the prior $p_{0}(\lambda)$ and the variational distribution $q(\lambda)$ are normally distributed, with means $(\mu_{0}=0,\mu_{q})$ and variances $(\sigma_{0}^{2}=1,\sigma_{q}^{2})$, the KL divergence can be written as:
$D_{KL}(q(\lambda)\|p_{0}(\lambda))=-\log{\sigma_{q}}+\frac{\sigma_{q}^{2}+\mu_{q}^{2}}{2}-\frac{1}{2}.$ (13)
2.2 Model architecture
Figure 1: DeepBayONet architecture.
$\mathcal{M}_{\theta}(t,x;\lambda)$ is a deep operator network (DeepONet) that consists of two components: the branch network and the trunk network [30, 31]. The branch network approximates the forward solution $y(t,x)$ of the PDE. The trunk network processes independent prior samples to generate samples of the approximate posterior $q(\lambda)$ that also influence the branch network’s predictions. See fig. 1 for a schematic representation of the network architecture. Our proposed architectue, hereafter referred to as DeepBayONet uses the DeepONet structure to estimate the forward solution and parameters of the PDE at the same time:
$y(t,x;\lambda)\approx\mathcal{M}_{\theta}(t,x,\lambda)=\sigma\Big{(}W_{0}\big{(}\mathsf{b}(t,x)\odot\mathsf{\tau}(\lambda)\big{)}+b_{0}\Big{)},$ (14)
where $W_{0}$ and $b_{0}$ are the output layer weights and bias, $\sigma(\cdot)$ is the activation function, and $\odot$ denotes the element-wise product. The branch network $\mathsf{b}(t,x)$ and trunk network $\mathsf{\tau}(\lambda)$ are general parametrized functions that can take the form of any neural network architectures including convolutional, recurrent, or transformer networks. The entire set of trainable parameters in the model is denoted by $\theta$.
Uncertainty estimation is critical in deep neural networks, particularly in scientific modeling and prediction tasks, to assess the reliability of predictions and the robustness of the model. There are two sources of uncertainty that can be considered: Aleatoric uncertainty arises from inherent noise in the observations, reflecting variability that cannot be reduced even with additional data. This type of uncertainty is often captured by modeling the variance of the output distribution, as demonstrated in probabilistic models such as heteroscedastic neural networks [32]. In contrast, epistemic uncertainty reflects the model’s lack of knowledge or suffiecient training about the underlying data distribution and can be reduced by incorporating additional data or improving the model architecture. Bayesian neural networks (BNNs) and dropout are among various methods that cab address epistemic uncertainty [32] by learning a distribution over the model parameters, allowing the model to express uncertainty in regions of the input space the model’s predictions are ambiguous. These two sources of uncertainty provide complementary information: aleatoric uncertainty quantifies noise in the data, while epistemic uncertainty assesses the model’s confidence in its predictions.
In the proposed model, both aleatoric and epistemic uncertainties are quantified explicitly through different components of the loss function and the model’s probabilistic formulation. Aleatoric uncertainty is captured through the observation noise variance $\sigma_{y}^{2}$, which governs the data likelihood term in the evidence lower bound (ELBO) formulation eq. 12 where $\sigma_{y}^{2}$ represents the variance of the observation noise. This term accounts for the inherent variability in the measurements, and by treating $\sigma_{y}^{2}$ as a learnable parameter, the model can adaptively estimate the aleatoric uncertainty directly from the data. High values of $\sigma_{y}^{2}$ indicate regions with significant observation noise, leading to reduced confidence in the predictions.
Epistemic uncertainty, on the other hand, is inferred from the posterior distribution $q(\lambda)$ over the latent parameters $\lambda$, as derived through variational inference. The posterior distribution is approximated by minimizing the Kullback-Leibler (KL) divergence in eq. 12. The variability in $q(\lambda)$, often characterized by its variance $\sigma_{q}^{2}$, quantifies the model’s uncertainty about $\lambda$. Furthermore, the residual noise variance $\sigma_{R}^{2}$, associated with the residuals of the governing equation eq. 8 also contributes to epistemic uncertainty by capturing how well the predicted solutions satisfy the physical constraints imposed by the partial differential equations (PDEs).
Through Monte Carlo sampling of eq. 12, the model evaluates both aleatoric and epistemic uncertainties simultaneously, enabling a robust and interpretable quantification of uncertainty in predictions.
2.3 Relation to Dropout
A distinctive feature of the DeepBayONet architecture is the interaction between the branch and trunk networks, which not only enables the model to represent the forward solution but also plays a role in uncertainty quantification. Specifically, the branch network $\mathsf{b}(t,x)$ approximates the forward operator by mapping the inputs $t$ and $x$ to a latent feature space. Simultaneously, the trunk network $\mathsf{\tau}(\lambda)$ processes samples of the latent parameter $\lambda$ drawn from the approximate posterior distribution $q(\lambda)$. These two networks interact through an element-wise product in eq. 14. The stochastic sampling of $\lambda\sim q(\lambda)$ introduces variability in $\mathsf{\tau}(\lambda)$, thereby perturbing the forward operator represented by $\mathsf{b}(t,x)$ during training.
This mechanism of perturbation bears a conceptual resemblance to dropout [33], widely-used regularization as well as epistemic (model) uncertainty quantification technique in neural networks. Dropout introduces noise during training by randomly setting activations to zero with a probability sampled from a Bernoulli distribution [33]. In DeepBayONet, instead of discrete binary perturbations, the trunk network introduces continuous perturbations drawn from the posterior distribution $q(\lambda)$. This continuous modulation of the branch network activations improves robustness while allowing the model to incorporate epistemic uncertainty in its predictions. Sampling from $q(\lambda)$ enables the model to account for uncertainty in the latent parameters $\lambda$. This uncertainty is propagated through the interaction between the branch and trunk networks. This continuous perturbation mechanism has two significant implications. First, it ensures that the model’s predictions generalize well by preventing the branch network from overfitting to deterministic solutions. Second, it provides a means to quantify epistemic uncertainty explicitly, as the variability in $\mathsf{\tau}(\lambda)$ represents the model’s confidence in its parameter estimates. Thus, the DeepONet architecture naturally integrates regularization and uncertainty quantification in a unified framework, leveraging the interaction between the branch and trunk networks to achieve both objectives.
2.4 Practical Aspects of the Loss Function
The loss function defined in eq. 12 enforces adherence to the governing equations of the dynamical system while capturing uncertainties in both the observations and the parameters. To ensure practical implementation, the total loss is decomposed into distinct components, each addressing specific aspects of the problem. Assuming $\mathcal{M}_{\theta}(t_{i},x_{i};\lambda)=\tilde{y}_{i}$ is the predicted PDE solution, the total loss is expressed as:
$\mathcal{L}=w_{\text{Interior}}L_{\text{Interior}}+w_{\text{IC}}L_{\text{IC}}+w_{\text{BC}}L_{\text{BC}}+w_{\text{STD}}L_{\text{STD}},$
where the individual loss components are defined to correspond with the probabilistic formulation in eq. 12.
To align these loss terms with the evidence lower bound (ELBO), variance-based scaling is introduced to reflect the uncertainties in the observations and residuals. The revised definitions of the individual components are:
$L_{\text{Interior}}(\theta,\lambda;X_{\text{Interior}})$ $=\frac{1}{|X_{\text{Interior}}|}\sum_{x\in\Omega}\frac{\left\|F(t,x,y;\lambda)\right\|^{2}}{\sigma_{R}^{2}},$ (15a)
$L_{\text{BC}}(\theta,\lambda;X_{\text{BC}})$ $=\frac{1}{|X_{\text{BC}}|}\sum_{x\in\partial\Omega}\frac{\left\|\mathcal{M}_{\theta}(t,x;\lambda)-y(t,x;\lambda)\right\|^{2}}{\sigma_{y}^{2}},$ (15b)
$L_{\text{IC}}(\theta,\lambda;X_{\text{IC}})$ $=\frac{1}{|X_{\text{IC}}|}\sum_{x\in\Omega}\frac{\left\|\mathcal{M}_{\theta}(0,x;\lambda)-y(0,x;\lambda)\right\|^{2}}{\sigma_{y}^{2}},$ (15c)
$L_{\text{STD}}(\theta,\lambda)$ $=\log{\sigma_{R}^{2}}+\log{\sigma_{y}^{2}}$ ) (15d)
where $\sigma_{R}^{2}$ is the variance of the residual noise and $\sigma_{y}^{2}$ is the variance of the observation noise. These variances play a critical role in balancing the contributions of the individual loss terms, ensuring that the model accounts for uncertainties appropriately.
The weights $w_{\text{Interior}},w_{\text{IC}},w_{\text{BC}},$ and $w_{\text{STD}}$ are hyperparameters used to balance the influence of the respective loss components. These weights can be selected through hyperparameter optimization or by employing adaptive weighting techniques, as explored in [34, 35]. The inclusion of variance scaling in eqs. 15a, 15b and 15c ensures consistency with the ELBO formulation, connecting the practical implementation to the probabilistic foundations of the model. This alignment also allows the model to accurately capture both aleatoric and epistemic uncertainties during training and inference.
3 Experiments
In this section, we train DeepBayONets on various algebraic and differential equations. In each case, the underlying equations have unknown parameters that are learned along with the forward dynamics. We compare the results with exact solutions and competing methodologies from the literature.
3.1 General Setup
The common configurations for all the experiments are as follows: each experiment utilizes a neural network architecture comprising three layers on both the branch and the trunk. The hyperbolic tangent ($\tanh$) activation function is employed universally. Training and optimization are conducted using the Adam optimizer. All implementations are performed in PyTorch, leveraging its advanced features for computational graphs and gradients. Specific details, such as the number of neurons per layer, epochs, batch size, and loss weights, are tuned for each experiment based on its requirements and complexity.
3.2 Forward Bayesian Regression Problem
This experiment demonstrates DeepBayONet’s capability to estimate aleatoric and epistemic uncertainty in a synthetic regression problem. The regression task is defined as follows:
$f(x)$ $=\sin\left(\frac{x}{2}\right),\quad$ $x\in[-1,1],$ (16)
$\epsilon_{i}$ $\sim\mathcal{N}\left(0,\sigma(x)^{2}\right),\quad$ $\sigma(x)^{2}=\frac{1-|x|}{16},$
$y_{i}$ $=f(x_{i})+\epsilon_{i},$
where $y_{i}$ are noisy observations with state-dependent inhomogeneous noise. The goal is to learn the function $f(x)$ while estimating two sources of uncertainty: Aleatoric uncertainty arising from state-dependent input noise in the data, and the epistemic uncertainty due to gaps in the data and testing outside the training range eq. 16.
The results are compared against various deep learning uncertainty quantification methods, including a Simple Neural Network with aleatoric uncertainty estimate (SNN), a Bayesian Neural Network (BNN), a Monte Carlo Dropout Network (MCDO), and a Deep Ensemble Network (DENN). Each architecture contains approximately 3,800–4,000 parameters to ensure a fair comparison. The data are split into a training set, an in-training-distribution (IDD) test set, and an out-of-training-distribution (OOD) test set.
The test problem and implementations are adopted from [36]. For a comprehensive review of uncertainty quantification methods in deep learning, refer to [37]. All networks are trained for 150 epochs with the Adam optimizer (batch size 16). Each model is run ten times with random initializations, and accuracies are averaged to mitigate random effects. Common hyperparameters are kept consistent across all models, while method-specific hyperparameters use default library values. To limit computational resource requirements, we have not done method-specific hyperparameter tuning.
We report training and testing mean squared errors (MSE) and the coverage of 95% confidence intervals (CI) predicted by the networks. This design facilitates an evaluation of both predictive accuracy and the quality of uncertainty estimates within and outside the training range.
Table 1: Mean and standard deviation of testing MSE over 10 runs
Architecture MSE (mean) MSE (std)
SNN 0.147 0.007
BNN 0.352 0.008
MCDO 0.198 0.034
DENN 0.154 0.012
DeepBayONet 0.143 0.005
Table 2: 95% confidence-interval coverage (averaged across 10 runs)
Architecture Total Testing IDD Coverage OOD Coverage
SNN 63.45 65.92 48.00
BNN 13.30 8.57 21.16
MCDO 52.15 41.35 72.49
DENN 59.70 36.21 88.16
DeepBayONet 89.05 85.42 97.00
The results demonstrate that DeepBayONet outperforms other architectures by achieving consistently lower mean squared errors (MSE) and significantly higher confidence-interval coverage. Notably, DeepBayONet excels in out-of-training-distribution scenarios, highlighting its robustness and reliability in quantifying uncertainty across diverse testing conditions.
3.3 One-dimensional function approximation
This experiment evaluates the ability of DeepBayONet to approximate a one-dimensional function while simultaneously estimating an unknown parameter and quantifying associated uncertainties. The function to be approximated is defined as:
$y(x)=\sin^{3}(\omega x),$ (17)
where $x\in[-1,1]$ and $\omega$ is an unknown parameter. For this experiment, the true value of $\omega=6$. The dataset consists of sampled values of $x$ from the domain paired with noisy observations:
$y^{\text{data}}(x)=y(x)+\varepsilon,\quad\varepsilon\sim\mathcal{N}(0,\sigma).$
Two noise scenarios are considered: a low-noise case with $\sigma=0.01$ and a high-noise case with $\sigma=0.1$.
Figure 2 shows the mean predicted function as well as the two standard deviation uncertainty bands for both high- and low-noise cases. The model demonstrates good agreement with the data within the training range, while predicting elevated uncertainty outside the training range. Higher observation noise ($\sigma=0.1$) results in more pronounced uncertainty, reflecting the model’s ability to adapt to noisy data.
The posterior distribution of the parameter $\omega$ is shown in fig. 3, with the mode close to the true value of $\omega=6.0$ in both cases. The narrower posterior distribution in the low-noise scenario ($\sigma=0.01$) indicates greater confidence in the parameter estimate compared to the high-noise scenario.
Figure 2: DeepBayONet learned solution for the regression problem in eq. 17. (a) $\sigma=0.1$. (b) $\sigma=0.01$
Figure 3: Posterior of the parameter $\omega$ learned by DeepBayONet from eq. 17.
3.4 One-dimensional unsteady heat equation
This experiment explores DeepBayONet’s capability to solve a one-dimensional unsteady heat equation while simultaneously estimating the diffusion parameter $D$ and the decay rate $\alpha$. The heat equation is given as:
[…]
where $D$ and $\alpha$ are unknown parameters. When $D=1$ and $\alpha=1$, the equation has the exact solution:
$y(t,x)=e^{-t}\sin(\pi x).$
A DeepBayONet model is trained to approximate this solution while learning the parameters $D$ and $\alpha$. The training dataset consists of 100 spatio-temporal samples with the corresponding exact solution $y(t_{i},x_{i})$. A standard normal prior is used for both parameters. Training is performed for 15,000 epochs with a batch size of 100, using the Adam optimizer and an initial learning rate of 0.01. The loss component weights are configured as follows:
$w_{\text{IC}}=3,\quad w_{\text{data}}=6,\quad w_{\text{BC}}=1,\quad w_{\text{int}}=1,\quad w_{\text{param std}}=1.$
The model contains 2,103 trainable parameters.
Figure 4: Loss components during the training of the 1D Heat equation.
During training, several metrics were recorded to monitor the model’s performance and convergence, including the total loss and individual loss components: PDE (Interior) loss, boundary condition (BC) loss, initial condition (IC) loss, data loss, and the standard deviation loss for predicted parameters (STD loss). These metrics are shown in fig. 4, highlighting the stability and convergence of the model.
Figure 5 compares a sample of the posterior predicted solution $y(t,x)$ with the exact solution. The grainy appearance of the predicted solution reflects the uncertainty in the parameters $D$ and $\alpha$, which affects the pointwise forward solution. However, obtaining multiple samples of the output and computing statistical measures, such as the mean and mode, is computationally efficient once the model is trained.
Figure 5: A sample of the predicted solution (left) and the exact solution (right) for the 1D Heat equation in eq. 18.
The histograms in fig. 6 illustrate the posterior distributions of the parameters $D$ and $\alpha$. The model successfully predicts these parameters, with the posterior mode aligning closely with the true values ($D=1.0$ and $\alpha=1.0$). Additionally, the secondary mode observed in the posterior distributions corresponds to larger values of $D$ and $\alpha$, indicating a highly diffusive and fast-decaying mode that still satisfies the PDE relatively well.
Figure 6: Histograms showing the distributions of Parameter $D$ and $\alpha$ for the 1D Heat equation in eq. 18.
Figure 7 further examines the errors and residuals for the DeepBayONet learned solution. Figure 7a shows the absolute error compared to the exact solution when using the mean of the posterior distribution of the parameters. Small errors are observed across the domain, except near the initial condition boundaries ($t=0,x=\pm 1$).
The PDE residuals, when evaluated using the primary mode (fig. 7b) and the secondary mode (fig. 7c) of the posterior distribution, provide additional insights. The secondary mode results in significantly larger residuals, reflecting suboptimal parameter values that deviate from the true solution.
Figure 7: Error and PDE residuals for the DeepBayONet learned solution of the unsteady heat eq. 18. (a) Forward solution absolute error compared to exact solution using the mean of the parameter posterior distribution. (b) PDE residual using the primary mode of the parameter posterior distribution. (c) PDE residual using the secondary mode of the posterior parameter distribution.
3.5 Two-dimensional Reaction-diffusion equation
This experiment evaluates DeepBayONet’s capability to solve a two-dimensional reaction-diffusion equation while estimating the reaction rate $k$. The equation is defined as:
$\lambda\left(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\right)+ku^{2}=f,\quad x,y\in[-1,1]^{2},$ (19a)
$f(x,y)=\frac{1}{50}-\pi^{2}u(x,y)+u(x,y)^{2},$ (19b)
$u(x,y)=\sin(\pi x)\sin(\pi y).$ (19c)
For $k=1$ and $\lambda=0.01$, the PDE has a known solution as given in eq. 19c. The training dataset consists of 100 randomly placed sensors within the domain that observe $u$ and $f$, with added observation noise of variance $0.01^{2}$. Additionally, 25 sensors are uniformly placed along each boundary to enforce Dirichlet boundary conditions.
Figure 8 shows the sensor positions along with contours of $u$ and $f$. The model is configured with 300 neurons per hidden layer, totaling 272,703 trainable parameters. It is trained for 35,000 epochs using the Adam optimizer with a batch size of 500. Loss component weights are calibrated as:
$w_{\text{Interior}}=20,000,\quad w_{\text{Data}}=60,000,\quad w_{\text{BC}}=100,\quad w_{\text{STD}}=20.$
This setup ensures the optimizer appropriately balances the various constraints.
(a) Training samples for $u$
(b) Training samples for $f$
Figure 8: Training samples, exact $u$ and $f$, and sensor positions for the 2D Reaction-diffusion equation in eq. 19.
Figure 9: Loss history plot for the 2D Reaction-diffusion equation in eq. 19. We observe a decrease in all loss components except for the variance of the posterior parameter distribution, as intentionally designed.
The training history, as shown in fig. 9, demonstrates consistent reduction in the PDE (Interior) and data losses during early epochs, followed by a decrease in the boundary condition loss. The variance of the posterior parameter distribution remains relatively unchanged due to the KL divergence term in the loss function.
Table 3 compares the estimated mean and standard deviation of the reaction rate $k$ using DeepBayONet and alternative Bayesian methods. DeepBayONet achieves results closest to the true value $k=1$ with the lowest standard deviation, demonstrating both accuracy and certainty in parameter estimation.
DeepBayONet B-PINN-HMC B-PINN-VI Dropout-1% Dropout-5%
Mean 0.999 1.003 0.895 1.050 1.168
Std $5.4\times 10^{-4}$ $5.75\times 10^{-3}$ $2.83\times 10^{-3}$ $2.00\times 10^{-3}$ $3.04\times 10^{-3}$
Table 3: 2D nonlinear diffusion-reaction system: Predicted mean and standard deviation for the reaction rate $k$ using different Bayesian PINN methods. The exact solution is $k=1$. Table reproduced from [29] with added data.
We conducted a second experiment by relaxing boundary condition enforcement ($w_{\text{BC}}=0$). Figure 10 shows the residuals of the PDE when using the mean posterior parameter value. Higher residuals are observed at the boundaries when boundary conditions are excluded from training.
(a) With boundary conditions
(b) Without boundary conditions
Figure 10: PDE residuals for the 2D Reaction-diffusion equation in eq. 19.
The posterior distributions of $k$ are shown in fig. 11. With boundary conditions, the posterior is narrow and centered around $k=1.0$. Without boundary conditions, the posterior is wider, skewed, and exhibits a long tail, indicating increased uncertainty towards positive reaction rates.
(a) With boundary conditions
(b) Without boundary conditions
Figure 11: Posterior distribution of the parameter $k$ for the 2D Reaction-diffusion equation in eq. 19.
Figure 12: Predicted mean solution, absolute error, and uncertainty for the 2D Reaction-diffusion equation in eq. 19 without boundary condition training.
Figure 12 shows the mean predicted solution and its error along with the predicted uncertainty for boundary-enforced training. In fig. 13, the corresponding plots for boundary-free training reveal reduced accuracy and elevated uncertainty, highlighting the ability of model to adapt to missing boundary conditions.
Figure 13: Predicted mean solution, absolute error, and uncertainty for the 2D Reaction-diffusion equation in eq. 19 with boundary condition training.
4 Conclusions
This study introduces Bayesian Deep Operator Networks (DeepBayONet) integrated with Physics-Informed Neural Networks (PINNs) to address parameter estimation and uncertainty quantification challenges in solving partial differential equations (PDEs). By employing a Bayesian framework with variational inference, the approach effectively captures both epistemic and aleatoric uncertainties, offering robust predictions in noisy and data-scarce environments. The method demonstrates versatility across tasks, including forward and inverse problems in one- and two-dimensional PDEs, with a relatively low-cost computational efficiency on par with maximum likelihood training as opposed to the traditional Bayesian neural network training.
The framework provides several advantages. It achieves comprehensive uncertainty quantification, enabling reliable predictions under complex scenarios, and maintains competitive accuracy. Additionally, the architecture integrates seamlessly with modern deep learning frameworks, allowing for scalability in addressing high-dimensional and nonlinear challenges.
However, the method has limitations. It requires careful tuning of hyperparameters, such as weights for loss components, which can introduce manual overhead. Although more efficient than many Bayesian alternatives, variational inference remains computationally intensive compared to deterministic methods.
Future work could address these challenges by exploring adaptive strategies for loss weighting to improve usability and generalizability. Further refinement of prior distributions to incorporate domain-specific knowledge could enhance parameter estimation and uncertainty quantification. We have focused on Gaussian priors in this work. Expanding the approach to complex parameter distributions and multiscale and multiphysics systems can broaden its applicability.
Acknowledgements
This research was supported by the CSML lab and the College of Engineering at California State University Long Beach. The work used Jetstream2 at Indiana University through allocation CIS230277 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.

### v2

@@
 By integrating data-driven learning with physical constraints, our method achieves robust and accurate solutions across diverse scenarios.
-Bayesian training is implemented through variational inference, allowing for comprehensive uncertainty quantification for both aleatoric and epistemic uncertainties.
+Bayesian training is implemented through variational inference, allowing for comprehensive uncertainty quantification for both data and model uncertainties.
 This ensures reliable predictions and parameter estimates even in noisy conditions or when some of the physical equations governing the problem are missing.
-The framework demonstrates its efficacy in solving forward and inverse problems, including the 1D unsteady heat equation and 2D reaction-diffusion equations, as well as regression tasks with sparse, noisy observations.
+The framework demonstrates its efficacy in solving forward and inverse problems, including the 1D unsteady heat equation, 2D reaction-diffusion equations, 3D eigenvalue problem, and various regression tasks with sparse, noisy observations.
 This approach provides a computationally efficient and generalizable method for addressing uncertainty quantification in PDE surrogate modeling.
@@
 Physics-Informed Neural Networks, Partial Differential Equations, Parameter Estimation, Machine Learning
-journal: arXiv
+journal: Computer Physics Communications
 […]
-[csulb]organization=Depratment of Computer Engineering and Computer Science, addressline= California State University, city=Long Beach, postcode=90840, state=CA, country=USA
+[csulb]organization=Department of Computer Engineering and Computer Science, addressline= California State University, city=Long Beach, postcode=90840, state=CA, country=USA
 1 Introduction
@@
 By embedding physical laws into the learning process, PINNs overcome the reliance on extensive datasets required by traditional neural networks, making them more practical for real-world scenarios [9].
-Applications of PINNs include solving fluid dynamics problems [10, 11], addressing inverse problems [12, 13], controlling dynamical systems [14, 15, 16], and performing uncertainty quantification [17].
+Applications of PINNs include solving fluid dynamics problems [10, 11], inverse problems [12, 13], controlling dynamical systems [14, 15, 16], and performing uncertainty quantification [17].
 Automatic differentiation (AD) is integral to the functionality of PINNs, enabling precise and efficient computation of PDE residuals, even for mixed and high-order derivatives.
@@
 [21] and Margossian [22].
-In scientific machine learning, probabilistic models are highly valued for estimating noise in observational data, addressing model errors, and accounting for incomplete knowledge of the underlying physics.
+In scientific machine learning, probabilistic models are highly valued for estimating noise in observational data, handling model errors, and accounting for incomplete knowledge of the underlying physics.
 Viana et al.
@@
 Li et al.
-[25] explore the challenges of solving inverse problems with PINNs, advocating for a Bayesian approach to provide robust uncertainty quantification, particularly when observations are sparse or noisy.
+[25] explore the challenges of solving inverse problems with PINNs, advocating a Bayesian approach to provide robust uncertainty quantification, particularly when observations are sparse or noisy.
 Bayesian Neural Networks (BNNs) [26, 27, 28] provide a probabilistic framework for uncertainty quantification by assigning distributions to network weights.
@@
 [29] introduced Bayesian Physics-Informed Neural Networks (B-PINNs), combining BNNs and PINNs to address forward and inverse PDE problems in noisy conditions.
-By integrating physical equations into training, B-PINNs offer a practical solution for probabilistic PDE learning tasks.
+By integrating physical equations into training, B-PINNs offer a practical solution to probabilistic PDE learning tasks.
 Accurately estimating PDE parameters while quantifying uncertainty remains a significant challenge, particularly in complex systems with scarce or noisy data.
+While Deep Operator Networks [30, 5, 8] can learn PDE solution operators from data, they provide deterministic predictions and lack a principled mechanism for uncertainty quantification.
+To resolve this limitation, B-DeepONet [31] extends the DeepONet framework by introducing a Bayesian formulation, enabling posterior sampling over the solution operator space.
+B-DeepONet learns model uncertainty by learning a posterior distribution over model parameters using Langevin dynamics.
+During training, a variant of Stochastic Gradient Langevin Dynamics (SGLD) applies gradients alternatively to the parameters of the branch and trunk networks to help reduce the computational cost of calculating gradients for the ensemble of parameters.
 The high computational demands and challenges in selecting suitable priors in Bayesian approaches restrict the practical use of black-box sampling methods in scientific machine learning.
 This presents an opportunity for new methodologies that effectively combine physics-guided training with computationally efficient Bayesian approaches.
-In this work, we propose a novel method that extends the capabilities of PINNs by integrating Bayesian methodologies with Deep Operator Networks (DeepONets) to enhance parameter estimation in PDEs.
-Our approach relies on variational inference: starting with a prior distribution for the unknown PDE parameters, we apply Bayesian inference to estimate the posterior distributions based on available data and physics.
+In this work, we propose a novel method that extends the capabilities of PINNs by integrating Bayesian methodologies with Deep Operator Networks (DeepONets) [32] to improve parameter estimation in PDEs.
+Our approach relies on variational inference: starting with a prior distribution for the unknown PDE parameters, we apply inference to estimate the posterior distribution of the PDE solution and the parameters based on the available data and known physics of the problem.
 The DeepOperator architecture modulates the forward prediction of the solution of the PDE with inverse parameter estimation.
-This method offers substantial computational advantages, providing uncertainty quantification comparable to B-PINNs but with significantly reduced computational costs.
-We continue our discussion in the following steps: In section 2, we detail the methodology of Bayesian Deep Operator Networks, demonstrate their application to parameter estimation in PDEs, and discuss their advantages and shortcomings compared to existing approaches.
-Section 3 presents experimental results on a variety of PDEs, showcasing the effectiveness of our method in solving forward and inverse problems.
+This method offers flexible uncertainty quantification and accurate differentiable solutions.
+We continue our discussion in the following steps: In section 2, we detail the methodology of Bayesian Deep Operator Networks, demonstrate their application to parameter estimation in PDEs, and discuss practical ways of quantifying model and data uncertainty in their predictions.
+Section 3 presents experimental results on a variety of PDEs, which showcase the effectiveness of our method in solving forward and inverse problems.
 Finally, we discuss the implications of our findings for future research and potential applications in various scientific and engineering domains in section 4.
+The source code for these experiments is publicly available as Jupyter Notebooks at https://github.com/csml-beach/deep-bayesian-operator-nets.
 2 Methodology
 2.1 Problem definition
-We start by considering a parametrized partial differential equation (PDE), defined over time span $t\in[t_{0},t_{f}]$ and spatial domain $x\in\Omega$:
+We start by considering a partial differential equation (PDE) with a latent parameter $\lambda$ defined over the time span $t\in[t_{0},t_{f}]$ and spatial domain $x\in\Omega$:
 ${F}(t,x,y;\lambda)=0,\quad(t,x)\in[t_{0},t_{f}]\times\Omega.$ (1)
@@
 1.
-We will start from a Bayesian formulation of this problem and derive a corresponding loss function for training the model.
-The posterior distribution over the latent parameter $\lambda$ given data $\mathcal{D}$ is expressed as
+We will start from a Bayesian formulation of this problem and derive a corresponding variational loss function for training the model.
+The posterior distribution over the latent parameter given the data is expressed as
 $p(\lambda\mid\mathcal{D})=\frac{p(\mathcal{D}\mid\lambda)\,p_{0}(\lambda)}{p(\mathcal{D})},$ (3)
 where $p_{0}(\lambda)$ represents the prior distribution over $\lambda$, $p(\mathcal{D}\mid\lambda)$ is the likelihood of the observed data given parameter $\lambda$, and $p(\mathcal{D})$ is the marginal likelihood or evidence.
-The likelihood term $p(\mathcal{D}\mid\lambda)$ is constructed from two components: the data likelihood $p(y_{i}\mid t_{i},x_{i},\lambda)$, which captures the relationship between the inputs $t_{i},x_{i}$ and outputs $y_{i}$, and the likelihood of the predicted solution satisfying the governing equations.
-The latter is achieved via the residuals of eq.
-1.
+The likelihood term $p(\mathcal{D}\mid\lambda)$ is constructed from two components: i) the data likelihood $p(y_{i}\mid t_{i},x_{i},\lambda)$ captures the relationship between the inputs $t_{i},x_{i}$ and outputs $y_{i}$.
+ii) and the likelihood of the predicted solution satisfying the governing equations.
+The latter is calculated based on the residuals of eq.
+1 when the model predicts the parameter and the solution.
+We refer to this residual as $\mathcal{R}_{\text{PDE}}$ hereafter.
 Assuming independence across training data points, the joint likelihood is written as
@@
 where $\sigma_{R}^{2}$ is the variance of the residual noise.
-$\mathcal{R}_{\text{PDE}}$ can either be simplified as $|F(t_{i},x_{i},y_{i},\lambda)|$ or modified to include other constraints such as the boundary and initial conditions.
+$\mathcal{R}_{\text{PDE}}$ can either be simplified as $\sum_{i}|F(t_{i},x_{i},y_{i},\lambda)|$ or modified to include other constraints such as the boundary and initial conditions.
 The corresponding log-likelihood of the residuals is
@@
 3 using variational inference (VI).
-Considering an approximate posterior distribution $q(\lambda)$ we minimize the Kullback-Leibler (KL) divergence:
+Considering an approximate posterior distribution $q(\lambda)$, we want to minimize the Kullback-Leibler (KL) divergence:
 $\operatorname*{arg\,min}_{q(\lambda)}D_{KL}(q(\lambda)\|p(\lambda\mid\mathcal{D}))=\int q(\lambda)\log\frac{q(\lambda)}{p(\lambda\mid\mathcal{D})}d\lambda.$ (9)
-Substituting eq. 3 for $p(\lambda\mid\mathcal{D})$ and rearranging terms, we obtain the well-known Evidence Lower Bound (ELBO) objective function:
+Substituting eq. 3 for $p(\lambda\mid\mathcal{D})$ and rearranging terms, yields the well-known Evidence Lower Bound (ELBO) objective function:
 […]
-Monte Carlo approximation of the integral in eq. 10 yields Expanding $p(\mathcal{D}\mid\lambda)$ using the joint likelihood, eq. 10 becomes
+Defining the expectation on $\lambda\sim q(\lambda)$ :
+[…]
+and separating $p(\mathcal{D}\mid\lambda)$ using the joint likelihood, eq. 10 becomes
 $\mathcal{L}_{\text{ELBO}}=\mathbb{E}_{q(\lambda)}\left[\sum_{i=1}^{N}\log p(y_{i}\mid x_{i},\lambda)+\log p(\mathcal{R}_{\text{PDE}}(\mathcal{M}_{\theta}(x_{i},\lambda)))\right]\\ -D_{KL}(q(\lambda)\|p_{0}(\lambda)).$ (11)
@@
 This enables a flexible uncertainty quantification method that can be utilized to estimate model or data uncertainty.
-We experiment with different variance estimation strategies in the experiments section 3.
+We provide examples for different uncertainty estimation strategies in the experiments section 3.
 The KL divergence term in eq.
-12 can be approximated empirically or analytically: If both the prior $p_{0}(\lambda)$ and the variational distribution $q(\lambda)$ are normally distributed, with means $(\mu_{0}=0,\mu_{q})$ and variances $(\sigma_{0}^{2}=1,\sigma_{q}^{2})$, the KL divergence can be written as:
+12 can be approximated empirically or analytically: If both the prior $p_{0}(\lambda)$ and the variational distribution $q(\lambda)$ are assumed to be normally distributed, with means $(\mu_{0}=0,\mu_{q})$ and variances $(\sigma_{0}^{2}=1,\sigma_{q}^{2})$, the KL divergence can be written as:
 $D_{KL}(q(\lambda)\|p_{0}(\lambda))=-\log{\sigma_{q}}+\frac{\sigma_{q}^{2}+\mu_{q}^{2}}{2}-\frac{1}{2}.$ (13)
 2.2 Model architecture
-Figure 1: DeepBayONet architecture.
-$\mathcal{M}_{\theta}(t,x;\lambda)$ is a deep operator network (DeepONet) that consists of two components: the branch network and the trunk network [30, 31].
+Figure 1: DeepBayONet feedforward architecture with trunk and branch networks. The trunk network learns the PDE parameters, while the branch network predicts the forward solution of the problem.
+We used a deep operator network (DeepONet) [33, 32] to parametrize $\mathcal{M}_{\theta}(t,x;\lambda)$.
+This architecture consists of two components: the branch network and the trunk network.
 The branch network approximates the forward solution $y(t,x)$ of the PDE.
-The trunk network processes independent prior samples to generate samples of the approximate posterior $q(\lambda)$ that also influence the branch network’s predictions.
+The trunk network processes independent prior samples from a standard normal distribution and transports them to samples of the approximate posterior $q(\lambda)$.
+The trunk and branch exchange information using an element-wise product before the output layers [30].
 See fig.
-1 for a schematic representation of the network architecture.
-Our proposed architectue, hereafter referred to as DeepBayONet uses the DeepONet structure to estimate the forward solution and parameters of the PDE at the same time:
+1 for a schematic representation of different components of the architecture.
+Our proposed network, referred to as DeepBayONet, uses the DeepONet architecture to estimate the forward solution and parameters of the PDE at the same time:
 $y(t,x;\lambda)\approx\mathcal{M}_{\theta}(t,x,\lambda)=\sigma\Big{(}W_{0}\big{(}\mathsf{b}(t,x)\odot\mathsf{\tau}(\lambda)\big{)}+b_{0}\Big{)},$ (14)
 where $W_{0}$ and $b_{0}$ are the output layer weights and bias, $\sigma(\cdot)$ is the activation function, and $\odot$ denotes the element-wise product.
-The branch network $\mathsf{b}(t,x)$ and trunk network $\mathsf{\tau}(\lambda)$ are general parametrized functions that can take the form of any neural network architectures including convolutional, recurrent, or transformer networks.
+The branch network $\mathsf{b}(t,x)$ and trunk network $\mathsf{\tau}(\lambda)$ are general parametrized functions and can be implemented using any neural network architectures, including convolutional, recurrent, or transformer networks.
 The entire set of trainable parameters in the model is denoted by $\theta$.
+2.3 DeepBayONet Uncertainty estimation
 Uncertainty estimation is critical in deep neural networks, particularly in scientific modeling and prediction tasks, to assess the reliability of predictions and the robustness of the model.
-There are two sources of uncertainty that can be considered: Aleatoric uncertainty arises from inherent noise in the observations, reflecting variability that cannot be reduced even with additional data.
-This type of uncertainty is often captured by modeling the variance of the output distribution, as demonstrated in probabilistic models such as heteroscedastic neural networks [32].
-In contrast, epistemic uncertainty reflects the model’s lack of knowledge or suffiecient training about the underlying data distribution and can be reduced by incorporating additional data or improving the model architecture.
-Bayesian neural networks (BNNs) and dropout are among various methods that cab address epistemic uncertainty [32] by learning a distribution over the model parameters, allowing the model to express uncertainty in regions of the input space the model’s predictions are ambiguous.
-These two sources of uncertainty provide complementary information: aleatoric uncertainty quantifies noise in the data, while epistemic uncertainty assesses the model’s confidence in its predictions.
-In the proposed model, both aleatoric and epistemic uncertainties are quantified explicitly through different components of the loss function and the model’s probabilistic formulation.
-Aleatoric uncertainty is captured through the observation noise variance $\sigma_{y}^{2}$, which governs the data likelihood term in the evidence lower bound (ELBO) formulation eq.
-12 where $\sigma_{y}^{2}$ represents the variance of the observation noise.
-This term accounts for the inherent variability in the measurements, and by treating $\sigma_{y}^{2}$ as a learnable parameter, the model can adaptively estimate the aleatoric uncertainty directly from the data.
+Two sources of uncertainty can be considered: data uncertainty arises from inherent noise in the observations, reflecting variability that cannot be reduced even with additional data.
+Under normal distribution assumptions, this type of uncertainty is often captured by modeling the observation error (co)variance.
+Specifically in deep learning, the predicted output of the network is augmented to include both the mean-prediction as well as the variance of the predicted output [34].
+In contrast, model uncertainty reflects the network’s lack of flexibility or insufficient training around the underlying data distribution and can be reduced by incorporating additional data or improving the model architecture.
+Bayesian neural networks (BNNs) and dropout are among various methods that can address epistemic uncertainty [34] by learning a distribution over the model parameters, allowing the model to express uncertainty in regions of the input space where the model’s predictions are ambiguous.
+These two sources of uncertainty provide complementary information: data uncertainty quantifies noise in the data, while model uncertainty assesses the model’s confidence in its predictions.
+In DeepBayONet we have the ability to quantify both sources of uncertainty.
+Data uncertainty can be captured through a learnable observation noise variance $\sigma_{y}^{2}(x;\theta)$ which appears in the data likelihood term of the ELBO loss eq.
+12 and accounts for the inherent variability in the measurements.
+By treating $\sigma_{y}^{2}$ as a learnable parameter, the model can adaptively estimate the data uncertainty based on the inputs.
 High values of $\sigma_{y}^{2}$ indicate regions with significant observation noise, leading to reduced confidence in the predictions.
-Epistemic uncertainty, on the other hand, is inferred from the posterior distribution $q(\lambda)$ over the latent parameters $\lambda$, as derived through variational inference.
-The posterior distribution is approximated by minimizing the Kullback-Leibler (KL) divergence in eq.
-12.
-The variability in $q(\lambda)$, often characterized by its variance $\sigma_{q}^{2}$, quantifies the model’s uncertainty about $\lambda$.
-Furthermore, the residual noise variance $\sigma_{R}^{2}$, associated with the residuals of the governing equation eq.
-8 also contributes to epistemic uncertainty by capturing how well the predicted solutions satisfy the physical constraints imposed by the partial differential equations (PDEs).
-Through Monte Carlo sampling of eq. 12, the model evaluates both aleatoric and epistemic uncertainties simultaneously, enabling a robust and interpretable quantification of uncertainty in predictions.
-2.3 Relation to Dropout
-A distinctive feature of the DeepBayONet architecture is the interaction between the branch and trunk networks, which not only enables the model to represent the forward solution but also plays a role in uncertainty quantification.
-Specifically, the branch network $\mathsf{b}(t,x)$ approximates the forward operator by mapping the inputs $t$ and $x$ to a latent feature space.
-Simultaneously, the trunk network $\mathsf{\tau}(\lambda)$ processes samples of the latent parameter $\lambda$ drawn from the approximate posterior distribution $q(\lambda)$.
-These two networks interact through an element-wise product in eq.
-14.
-The stochastic sampling of $\lambda\sim q(\lambda)$ introduces variability in $\mathsf{\tau}(\lambda)$, thereby perturbing the forward operator represented by $\mathsf{b}(t,x)$ during training.
-This mechanism of perturbation bears a conceptual resemblance to dropout [33], widely-used regularization as well as epistemic (model) uncertainty quantification technique in neural networks.
+Model uncertainty in DeepBayONet is inferred by randomly perturbing the latent representation $h(t,x;\lambda)=b(t,x;\lambda)$ of the forward solution.
+This latent representation is taken from the deep layers in the branch network before the output layer (see fig.
+1).
+This mechanism of perturbation bears a conceptual resemblance to dropout [35], widely used for regularization as well as epistemic (model) uncertainty quantification.
 Dropout introduces noise during training by randomly setting activations to zero with a probability sampled from a Bernoulli distribution [35].
-In DeepBayONet, instead of discrete binary perturbations, the trunk network introduces continuous perturbations drawn from the posterior distribution $q(\lambda)$.
-This continuous modulation of the branch network activations improves robustness while allowing the model to incorporate epistemic uncertainty in its predictions.
-Sampling from $q(\lambda)$ enables the model to account for uncertainty in the latent parameters $\lambda$.
-This uncertainty is propagated through the interaction between the branch and trunk networks.
+Let ${h}\in\mathbb{R}^{d}$ denote the hidden activations of a layer.
+During dropout, each element $h_{i}$ is multiplied by a binary random variable $z_{i}$ drawn independently from a Bernoulli distribution:
+$z_{i}$ $\sim\mathrm{Bernoulli}(p),\quad\text{for }i=1,\dots,d,$ (15a)
+$\tilde{{h}}$ $={z}\odot{h},$ (15b)
+[…]
+Our framework adopts a similar stochastic perturbation principle to that of dropout, but with a key difference: DeepBayONet introduces continuous perturbations to the solution representation.
+These perturbations are drawn from a distribution over perturbations learned by the trunk network:
+$\lambda$ $\sim\mathcal{N}(0,\mathbf{I}),$ (16a)
+$\tilde{{h}}$ $=\tau(\lambda)\odot{h}.$ (16b)
 This continuous perturbation mechanism has two significant implications.
 First, it ensures that the model’s predictions generalize well by preventing the branch network from overfitting to deterministic solutions.
-Second, it provides a means to quantify epistemic uncertainty explicitly, as the variability in $\mathsf{\tau}(\lambda)$ represents the model’s confidence in its parameter estimates.
-Thus, the DeepONet architecture naturally integrates regularization and uncertainty quantification in a unified framework, leveraging the interaction between the branch and trunk networks to achieve both objectives.
+Second, it provides a means to quantify the discovered parameters’ uncertainty, as the variability in $\mathsf{\tau}(\lambda)$ represents the model’s confidence in its parameter estimates.
+The posterior distribution $q(\lambda)$ over the latent parameters can be empirically estimated using the trunk network to understand the model’s confidence in the predicted parameter.
+Finally, we note that if one is interested in estimating the residual noise variance $\sigma_{R}^{2}$, associated with the residuals of the governing equation eq.
+8 It can also be learned as a parametric function of the network.
+The DeepBayONet architecture naturally integrates regularization and uncertainty quantification in a unified framework, leveraging the interaction between the branch and trunk networks to achieve both objectives.
 2.4 Practical Aspects of the Loss Function
-The loss function defined in eq.
-12 enforces adherence to the governing equations of the dynamical system while capturing uncertainties in both the observations and the parameters.
-To ensure practical implementation, the total loss is decomposed into distinct components, each addressing specific aspects of the problem.
+It is possible to numerically optimize the cost function eq.
+12 through Monte Carlo sampling over $t,x,\lambda$.
+When input samples are drawn from the training data, the data likelihood term will guide the model to predict the known forward solution.
+When inputs are not in the dataset, the residual likelihood trains the model to adhere to physical constraints and the PDE equations.
+Simultaneously, the entropy term in eq.
+12 encourages the posterior of the learned parameter to remain normally distributed.
+A common practice in the PINNs literature is to scale different components of the loss function to give importance to certain aspects and de-emphasize others.
+We follow the same practical technique and assign different weights to the terms in eq.
+12.
 Assuming $\mathcal{M}_{\theta}(t_{i},x_{i};\lambda)=\tilde{y}_{i}$ is the predicted PDE solution, the total loss is expressed as:
-$\mathcal{L}=w_{\text{Interior}}L_{\text{Interior}}+w_{\text{IC}}L_{\text{IC}}+w_{\text{BC}}L_{\text{BC}}+w_{\text{STD}}L_{\text{STD}},$
-where the individual loss components are defined to correspond with the probabilistic formulation in eq. 12.
-To align these loss terms with the evidence lower bound (ELBO), variance-based scaling is introduced to reflect the uncertainties in the observations and residuals.
-The revised definitions of the individual components are:
+$\mathcal{L}=w_{\text{Interior}}L_{\text{Interior}}+w_{\text{IC}}L_{\text{IC}}+w_{\text{BC}}L_{\text{BC}}+w_{\text{STD}}L_{\text{STD}}+w_{\text{Data}}L_{\text{Data}},$ (17)
+where
 $L_{\text{Interior}}(\theta,\lambda;X_{\text{Interior}})$ $=\frac{1}{|X_{\text{Interior}}|}\sum_{x\in\Omega}\frac{\left\|F(t,x,y;\lambda)\right\|^{2}}{\sigma_{R}^{2}},$ (18a)
@@
 $L_{\text{IC}}(\theta,\lambda;X_{\text{IC}})$ $=\frac{1}{|X_{\text{IC}}|}\sum_{x\in\Omega}\frac{\left\|\mathcal{M}_{\theta}(0,x;\lambda)-y(0,x;\lambda)\right\|^{2}}{\sigma_{y}^{2}},$ (18c)
-$L_{\text{STD}}(\theta,\lambda)$ $=\log{\sigma_{R}^{2}}+\log{\sigma_{y}^{2}}$ ) (15d)
-where $\sigma_{R}^{2}$ is the variance of the residual noise and $\sigma_{y}^{2}$ is the variance of the observation noise.
-These variances play a critical role in balancing the contributions of the individual loss terms, ensuring that the model accounts for uncertainties appropriately.
-The weights $w_{\text{Interior}},w_{\text{IC}},w_{\text{BC}},$ and $w_{\text{STD}}$ are hyperparameters used to balance the influence of the respective loss components.
-These weights can be selected through hyperparameter optimization or by employing adaptive weighting techniques, as explored in [34, 35].
-The inclusion of variance scaling in eqs.
-15a, 15b and 15c ensures consistency with the ELBO formulation, connecting the practical implementation to the probabilistic foundations of the model.
-This alignment also allows the model to accurately capture both aleatoric and epistemic uncertainties during training and inference.
+$L_{\text{Data}}(\theta,\lambda;X_{\text{Data}})$ $=\frac{1}{|X_{\text{Data}}|}\sum_{(x)\in X_{\text{Data}}}\frac{\left\|\mathcal{M}_{\theta}(t,x;\lambda)-y(t,x)\right\|^{2}}{\sigma_{y}^{2}},$ (18d)
+$L_{\text{STD}}(\theta,\lambda)$ $=\left\|\log{\sigma_{y}^{2}}\right\|,$ (18e)
+where we have expanded the log-likelihood of the data term in eq.
+12 to also include any boundary and initial conditions we may want to enforce for the learned solution.
+[^1] The loss component in eq.
+18e was found empirically to be simpler than evaluating the exact KL-divergence.
+The intuition behind it is that the log of the variance of the predicted solution needs to be of moderate magnitude.
+Too large and too small variances are discouraged.
+The weights $w_{\text{Interior}},w_{\text{IC}}$, etc., are inevitably hyperparameters that balance the influence of the respective loss components, and their choice can highly affect the quality of the learned solution.
+Adaptively choosing these weights is an active area of research in PINNs [36, 37].
+In this paper, we have used a combination of grid search and hand-tuning of hyperparameters.
 3 Experiments
 In this section, we train DeepBayONets on various algebraic and differential equations.
-In each case, the underlying equations have unknown parameters that are learned along with the forward dynamics.
-We compare the results with exact solutions and competing methodologies from the literature.
-3.1 General Setup
-The common configurations for all the experiments are as follows: each experiment utilizes a neural network architecture comprising three layers on both the branch and the trunk.
+In many cases, the underlying equations have unknown parameters that are learned along with the forward dynamics.
+We compare the results with exact solutions and competing methodologies from the literature when they are available.
+3.1 General experiment setup
+The common configurations used for all the experiments are as follows: each experiment utilizes a feedforward neural network architecture comprising three layers on both the branch and the trunk.
 The hyperbolic tangent ($\tanh$) activation function is employed universally.
 Training and optimization are conducted using the Adam optimizer.
-All implementations are performed in PyTorch, leveraging its advanced features for computational graphs and gradients.
+All implementations use the PyTorch package, leveraging its features for computational graphs and gradients.
 Specific details, such as the number of neurons per layer, epochs, batch size, and loss weights, are tuned for each experiment based on its requirements and complexity.
+These experiments are documented in the paper’s GitHub repository [^2] https://github.com/csml-beach/deep-bayesian-operator-nets for reference.
 3.2 Forward Bayesian Regression Problem
-This experiment demonstrates DeepBayONet’s capability to estimate aleatoric and epistemic uncertainty in a synthetic regression problem. The regression task is defined as follows:
+This experiment demonstrates DeepBayONet’s capability to handle uncertainties in a synthetic regression problem. The regression task is defined as follows:
 $f(x)$ $=\sin\left(\frac{x}{2}\right),\quad$ $x\in[-1,1],$ (19)
@@
 where $y_{i}$ are noisy observations with state-dependent inhomogeneous noise.
-The goal is to learn the function $f(x)$ while estimating two sources of uncertainty: Aleatoric uncertainty arising from state-dependent input noise in the data, and the epistemic uncertainty due to gaps in the data and testing outside the training range eq.
-16.
+The goal is to learn the function $f(x)$ while estimating two sources of uncertainty: data(Aleatoric) uncertainty arising from state-dependent input noise in the data, and the model(epistemic) uncertainty due to the gap between the training data and inference points.
+Figure 2 shoes the dataset for this problem.
 The results are compared against various deep learning uncertainty quantification methods, including a Simple Neural Network with aleatoric uncertainty estimate (SNN), a Bayesian Neural Network (BNN), a Monte Carlo Dropout Network (MCDO), and a Deep Ensemble Network (DENN).
@@
 The data are split into a training set, an in-training-distribution (IDD) test set, and an out-of-training-distribution (OOD) test set.
-The test problem and implementations are adopted from [36].
+The latter division is used to gauge the predicted uncertainty of the model outside the training range: Ideally, the model would become less confident the further away the inference point is from the training data.
+In fig.
+2 the training range is denoted with vertical blue lines.
+This test problem and the suite of uncertainty-aware deep neural networks are adopted from [38].
 For a comprehensive review of uncertainty quantification methods in deep learning, refer to [39].
@@
 Common hyperparameters are kept consistent across all models, while method-specific hyperparameters use default library values.
-To limit computational resource requirements, we have not done method-specific hyperparameter tuning.
+To limit computational resource requirements, we have not performed method-specific hyperparameter tuning.
 We report training and testing mean squared errors (MSE) and the coverage of 95% confidence intervals (CI) predicted by the networks.
@@
 Notably, DeepBayONet excels in out-of-training-distribution scenarios, highlighting its robustness and reliability in quantifying uncertainty across diverse testing conditions.
+Figure 2: (a) DeepBayONet without trunk (SNN).
+(b) Monte Carlo Dropout (MCDO).
+(c) Deep Bayesian Operator Network (DeepBayONet).
+Each method exhibits distinct uncertainty quantification behavior under the same experimental conditions.
+The two vertical blue lines in each plot denote the division between the IDD dataset (between the lines) and ODD dataset (everywhere else).
+Figure 2 presents the uncertainty quantification for DeepBayONet without trunk (SNN), Monte Carlo Dropout Network, and regular DeepBayONet.
+SNN effectively predicts the uncertainty in the data within the training range accurately; however, it estimates relatively low uncertainty for data outside the training range.
+This behavior is attributed to the lack of any proper mechanism to quantify the uncertainty in the network parameters.
+On the other hand, MCDO predicts a higher uncertainty outside of the training range.
+This results from the activation perturbations induced by the dropout.
+Yet, MCDO exhibits overconfident prediction on inference points within the training range.
+Compared to figure 2a, which depicts DeepBayONet without the trunk, the regular DeepBayONet effectively captures meaningful uncertainty by perturbing the activation of the intermediate layer of the network.
+As a result, DeepBayONet accurately captures both sources of uncertainty for both data within and outside the training range as indicated in figure 2c.
 3.3 One-dimensional function approximation
@@
 The posterior distribution of the parameter $\omega$ is shown in fig.
-3, with the mode close to the true value of $\omega=6.0$ in both cases.
+4, with the mode close to the true value of $\omega=6.0$ in both cases.
 The narrower posterior distribution in the low-noise scenario ($\sigma=0.01$) indicates greater confidence in the parameter estimate compared to the high-noise scenario.
@@
 The loss component weights are configured as follows:
-$w_{\text{IC}}=3,\quad w_{\text{data}}=6,\quad w_{\text{BC}}=1,\quad w_{\text{int}}=1,\quad w_{\text{param std}}=1.$
+$w_{\text{IC}}=3,\quad w_{\text{Data}}=6,\quad w_{\text{BC}}=1,\quad w_{\text{Interior}}=1,\quad w_{\text{STD}}=1.$
 The model contains 2,103 trainable parameters.
-Figure 4: Loss components during the training of the 1D Heat equation.
+A basic grid search was employed to navigate the vast hyperparameter space and ensure the reproducibility of the experiments.
+Additionally, we implemented a learning rate scheduler that adaptively reduces the learning rate, further minimizing the model’s sensitivity to hyperparameters and improving its ability to navigate local minima.
+It’s worth noting that manually examining the magnitude of each component of the loss function eq.
+17 between training epochs is also an option to adjust them proportionally.
 During training, several metrics were recorded to monitor the model’s performance and convergence, including the total loss and individual loss components: PDE (Interior) loss, boundary condition (BC) loss, initial condition (IC) loss, data loss, and the standard deviation loss for predicted parameters (STD loss).
 These metrics are shown in fig.
-4, highlighting the stability and convergence of the model.
+5a, highlighting the stability and convergence of the model.
+Figure 5: (a) Loss history plot for 1D Heat equation in eq. 21 eq. 21. (b) Loss history plot for the 2D Reaction-diffusion equation in eq. 22. (c) Loss history plot for 3D Helmholtz equation in eq. 23
 Figure 6 compares a sample of the posterior predicted solution $y(t,x)$ with the exact solution.
-The grainy appearance of the predicted solution reflects the uncertainty in the parameters $D$ and $\alpha$, which affects the pointwise forward solution.
+The grainy appearance of the predicted solution reflects the uncertainty in the parameters $D$ and $\alpha$, which affects the point-wise forward solution.
 However, obtaining multiple samples of the output and computing statistical measures, such as the mean and mode, is computationally efficient once the model is trained.
@@
 Small errors are observed across the domain, except near the initial condition boundaries ($t=0,x=\pm 1$).
-The PDE residuals, when evaluated using the primary mode (fig.
-7b) and the secondary mode (fig.
-7c) of the posterior distribution, provide additional insights.
-The secondary mode results in significantly larger residuals, reflecting suboptimal parameter values that deviate from the true solution.
 Figure 8: Error and PDE residuals for the DeepBayONet learned solution of the unsteady heat eq.
-18.
+21.
 (a) Forward solution absolute error compared to exact solution using the mean of the parameter posterior distribution.
@@
 (c) PDE residual using the secondary mode of the posterior parameter distribution.
+The PDE residuals, when evaluated using the primary mode (fig.
+8b) and the secondary mode (fig.
+8c) of the posterior distribution, provide additional insights.
+The secondary mode results in significantly larger residuals, reflecting suboptimal parameter values that deviate from the true solution.
 3.5 Two-dimensional Reaction-diffusion equation
@@
 For $k=1$ and $\lambda=0.01$, the PDE has a known solution as given in eq.
-19c.
+22c.
 The training dataset consists of 100 randomly placed sensors within the domain that observe $u$ and $f$, with added observation noise of variance $0.01^{2}$.
@@
 Figure 9: Training samples, exact $u$ and $f$, and sensor positions for the 2D Reaction-diffusion equation in eq. 22.
-Figure 9: Loss history plot for the 2D Reaction-diffusion equation in eq.
-19.
-We observe a decrease in all loss components except for the variance of the posterior parameter distribution, as intentionally designed.
-The training history, as shown in fig.
-9, demonstrates consistent reduction in the PDE (Interior) and data losses during early epochs, followed by a decrease in the boundary condition loss.
-The variance of the posterior parameter distribution remains relatively unchanged due to the KL divergence term in the loss function.
+Comparing the training loss history, as shown in fig.
+5, we observe that the reaction-diffusion losses in fig.
+5b have magnitudes of difference, while in fig.
+5a the loss components are well-balanced.
+This is due to the fact that in fig.
+5a the weights for different losses are about the same while the data loss and interior loss for the reaction-diffusion equation had to be scaled up dramatically for a successful training.
 Table 3 compares the estimated mean and standard deviation of the reaction rate $k$ using DeepBayONet and alternative Bayesian methods.
@@
 In fig.
-13, the corresponding plots for boundary-free training reveal reduced accuracy and elevated uncertainty, highlighting the ability of model to adapt to missing boundary conditions.
+13, the corresponding plots for boundary-free training reveal reduced accuracy and elevated uncertainty, highlighting the ability of the model to adapt to missing boundary conditions.
 Figure 13: Predicted mean solution, absolute error, and uncertainty for the 2D Reaction-diffusion equation in eq. 22 with boundary condition training.
+3.6 3D Eigenvalue problem in a unit ball
+In this experiment, we use the three–dimensional Dirichlet–Laplacian 3D problem:
+$\begin{cases}-\Delta u(\mathbf{x})=\lambda\,u(\mathbf{x}),&\|\mathbf{x}\|<1,\\[4.0pt] u(\mathbf{x})=0,&\|\mathbf{x}\|=1.\end{cases}$ (23)
+In Cartesian coordinates the domain boundary $x^{2}+y^{2}+z^{2}=1$ cuts through a regular grid at oblique angles.
+Conventional grid-based solvers must either deform a Cartesian mesh to fit the curved boundary or use immersed-boundary tricks.
+Both options complicate stencil design.
+Furthermore, estimating the eigenvalue $\lambda$ as a parameter depends on the entire error in $u$, not just local residuals, which makes accurate estimation of both quantities challenging.
+We solve this problem using DeepBayONet by approximating the eigenmode via the branch network and the eigenvalue using the trunk network.
+A numerical challenge in learning the eigenpairs is that the PDE and the boundary conditions are satisfied by $u\equiv 0$ for any $\lambda$, so additional constraints are required to force the network toward a non‐trivial eigenmode.
+$\mathcal{L}=\underbrace{\mathbb{E}_{\Omega}\!\bigl{[}-\Delta u_{\theta}-\lambda_{\theta}u_{\theta}\bigr{]}^{2}}_{\text{PDE residual}}\;+\;\underbrace{\mathbb{E}_{\partial\Omega}\!\bigl{[}u_{\theta}^{2}\bigr{]}}_{\text{Dirichlet BC}}\;+\;\underbrace{\bigl{(}\mathbb{E}_{\Omega}[u_{\theta}^{2}]-1\bigr{)}^{2}}_{\text{normalization}}.$
+The loss function for this problem combines three terms: (i) the PDE residual $(-\Delta u-\lambda u)^{2}$ evaluated at random interior points, (ii) the squared boundary condition on the sphere, and (iii) a simple normalization $(\mathbb{E}[u^{2}]-1)^{2}$ to guide the learned solution away from the trivial solution $u\equiv 0$.
+Automatic differentiation supplies exact gradients of all three terms, so the optimizer can update the network weights and the spectral parameter $\lambda$ in a single pass.
+The eigen-functions of eq.
+23 are available in closed form.
+They separate into spherical Bessel functions $j_{\ell}(k_{\ell s}r)$ and spherical harmonics $Y_{\ell}^{m}(\theta,\varphi)$.
+Our neural solver never sees this analytic structure.
+Throughout the experiments, the network receives only Cartesian coordinates $(x,y,z)$ and must satisfy the boundary condition $x^{2}+y^{2}+z^{2}=1$ implicitly through the loss.
+Working in Cartesian space, therefore, removes the coordinate system in which the PDE decouples, forces the optimizer to discover a non-trivial eigenfunction from scratch, and provides a more demanding test of the physics-informed loss.
+The first two eigenvalues are $\lambda_{01}=\pi^{2}\approx 9.87$ (radial mode) and $\lambda_{11}=4.49341^{2}\approx 20.19$ (dipole mode).
+During training, we monitor convergence toward $\lambda_{11}$.
+Figure 14 presents a comprehensive visualization of the model’s performance on the 3D Helmholtz eigenvalue problem using both equatorial slices at fixed values and a meridional slice at $x=0$ In each slice, the model’s predicted field (left), the exact solution (center), and the point-wise absolute error (right) are shown.
+The polar plots for $z=0.0,0.25$ and $z=0.50$ illustrate that the model learns the correct dipole structure expected.
+The meridional slice, shown in the bottom row, further verifies the agreement between predicted and exact solutions in the full 3D domain.
+The errors are highly localized and remain within a small magnitude across the domain.
+To further assess the model’s accuracy along a 1D cut, Figure 15b shows the predicted and exact values of the solution along the central vertical axis $(x=y=0).$
+Complementing this, the Figure 15a histogram of the learned eigenvalue parameter $\lambda$ confirms sharp convergence to the ground truth value $\lambda\approx 4.493$, validating that the network accurately identifies the underlying spectral structure.
+Predicted
+Exact
+Error
+$z=0.00$
+$z=0.25$
+$z=0.50$
+$x=0$
+Figure 14: Combined results for the 3D Helmholtz experiment.
+(a) 3D Axis Plot
+(b) 3D Histogram
+Figure 15: Axis and Histogram plot for the 3D Helmholtz equation in eq. 23
 4 Conclusions
 This study introduces Bayesian Deep Operator Networks (DeepBayONet) integrated with Physics-Informed Neural Networks (PINNs) to address parameter estimation and uncertainty quantification challenges in solving partial differential equations (PDEs).
-By employing a Bayesian framework with variational inference, the approach effectively captures both epistemic and aleatoric uncertainties, offering robust predictions in noisy and data-scarce environments.
-The method demonstrates versatility across tasks, including forward and inverse problems in one- and two-dimensional PDEs, with a relatively low-cost computational efficiency on par with maximum likelihood training as opposed to the traditional Bayesian neural network training.
+By employing a Bayesian framework with variational inference, the approach effectively captures both data and model uncertainties, offering robust predictions in noisy and data-scarce environments.
+The method demonstrates versatility across tasks, including regression and PDE problems in one, two, and three dimensions.
 The framework provides several advantages.
 It achieves comprehensive uncertainty quantification, enabling reliable predictions under complex scenarios, and maintains competitive accuracy.
-Additionally, the architecture integrates seamlessly with modern deep learning frameworks, allowing for scalability in addressing high-dimensional and nonlinear challenges.
-However, the method has limitations.
-It requires careful tuning of hyperparameters, such as weights for loss components, which can introduce manual overhead.
-Although more efficient than many Bayesian alternatives, variational inference remains computationally intensive compared to deterministic methods.
-Future work could address these challenges by exploring adaptive strategies for loss weighting to improve usability and generalizability.
-Further refinement of prior distributions to incorporate domain-specific knowledge could enhance parameter estimation and uncertainty quantification.
-We have focused on Gaussian priors in this work.
-Expanding the approach to complex parameter distributions and multiscale and multiphysics systems can broaden its applicability.
-Acknowledgements
+Additionally, the architecture integrates seamlessly with modern deep learning frameworks, allowing GPU-accelerated applications.
+The method has limitations that may be addressed in future directions: Careful tuning of hyperparameter can be achieved algorithmically by adaptive strategies for loss weighting.
+Further refinements that can use non-Gaussian priors to incorporate domain-specific knowledge could enhance parameter estimation and uncertainty quantification.
+Finally, although DeepBayONet discovers the parameters of the PDE from the available data and physics, it does not learn a parametrized solution operator of the PDE.
+We leave that extension to future endeavors.
+Acknowledgments
+We thank the reviewers for their careful reading and thoughtful feedback, which helped us improve this work.
 This research was supported by the CSML lab and the College of Engineering at California State University Long Beach.